OpenStudy (clamin):

PLEASE HELP!! In a sample of 400 lithium atoms, it is found that 30 atoms are lithium-6 and 370 atoms are lithium-7. Calculate the average atomic mass of lithium.

OpenStudy (anonymous):

isotopic mass of lithium 6 is 6.015, and for 7 is 7.016(Wikipedia). now % of abundance of 6 isomer is equal to 30/400=7.5%, % of abundance of 7 isomer is equal to 370/400=92.5% so av atomic mass= ((mass of the 1st isotope * relative abundance )+(mass of 2nd isotope*relative abundance)) so av mass =((6.015*0.075)+(7.016*.925) =6.9409